OpenStudy (anonymous):

Sarah took the advertising deparment from her company on a round trip to chicago to meet with a potential client. Including Sarah, a total of 14 people took the trip. She was able to purchase coach tickets for $250 and first class tickets for $1150. She used her total budget for airfare for the trip, which was $10,700. How many first class tickets did she buy? How many coach tickets did she buy?

OpenStudy (anonymous):

250x+1150y=10,700 x+y=14 so y=14-x 250x+1150(14-x)=10700